OpenStudy (jennychan12):

Find the points of any horizontal tangent lines: (x^2+xy+y^2) = 6. Am I doing something wrong?

OpenStudy (jennychan12):

\[x^2+xy +y^2 = 6 \] \[2x+(x+y)y'+2yy'=0\] \[y'(x+y+2y) = -2x\] \[y'=\frac{ -2x }{ x+3y }\]

OpenStudy (jennychan12):

I'm thinking set y' = 0. Am I correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

check the derivative of your middle term: [xy]' = ???

OpenStudy (jennychan12):

x+y ? product rule...?

OpenStudy (jennychan12):

ohhhh i seeeeeeeee

OpenStudy (anonymous):

yes... product rule: \(\large [xy]'=x'y+y'x=y+y'x \)

OpenStudy (jennychan12):

\[2x+xy'+y+2yy' = 0\] \[y'(x+2y) = -2x-y\] \[y' = \frac{ -2x-y }{ x+2y }\] so then \[-2x = y \] ?

OpenStudy (anonymous):

yep... now you need to find WHERE on the graph this occurs: y = -2x so substitute this into the original equation and solve...

OpenStudy (anonymous):

\(\large x^2+xy+y^2=6\), \(\large y=-2x \) \(\large x^2+x(-2x)+(-2x)^2=6\) ????

OpenStudy (jennychan12):

ohhh i remember this now. sorry i learned this quite a while ago, and forgot some parts :'(

OpenStudy (anonymous):

ok... cya....

OpenStudy (jennychan12):

thanks!